What volume of 95 mass% sulphuric acidand(denisty=1.85g/cm³) what mass of water must be taken to prepare 100cm³ of 15 mass% solution sulphuirc acid(density=1.10g/cm³)?

Grade:12

2 Answers

Arun
5 years ago
15% solution
Mass = volume* density
volume=100 cm3
density=1.10 g/cm3
mass of 15% solution = 100*1.10=110 g
mass of sulphuric acid in it =.15*110=16.5g
mass of water present in it =110-16.5=93.5 g of water
 95% solution
let mass of 95% solution required to prepare 15% H2SO4 with 16.5 g acid in it be x
16.5g =0.95* x
x=16.5*100/95 =17.4 g
Volume of 95%H2SO4 needed to prepare 15%H2SO4
VOLUME =mass/density
Volume =17.4/1.85= 9.4 cm3
Volume water needed to make up the volume =100-9.4=90.6 cm3
 
Kushagra Madhukar
3 years ago
Dear student,
Please find the attached answer to your problem.
 
15% solution
Mass = volume* density
volume=100 cm3
density=1.10 g/cm3
mass of 15% solution = 100*1.10=110 g
mass of sulphuric acid in it =.15*110=16.5g
mass of water present in it =110-16.5=93.5 g of water 
95% solution
let mass of 95% solution required to prepare 15% H2SO4 with 16.5 g acid in it be x
16.5g =0.95* x
x=16.5*100/95 =17.4 g
Volume of 95%H2SO4 needed to prepare 15%H2SO4
VOLUME =mass/density
Volume =17.4/1.85= 9.4 cm3
Volume water needed to make up the volume =100-9.4=90.6 cm3
 
Regards,
Kushagra